Dublin University Rifle Club

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Mr. R.R.Green 29 Magherana Park Waringstown Craigavon Co.Armagh BT66 7QF

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Air Rifle Round 3  18th February 2002

A Chara,

Enclosed are the air rifle cards for round 3. My apologies for the slight delay Robert, but I’ve been somewhat ill for the past fortnight (it’s ’flu season down here).

Misé le meas,

Mark Dennehy, DURC Secretary